# On generating mapping class groups by pseudo-Anosov elements

### v2

@@ 2. Preliminaries
First, we show the former part of Lemma 3.4 (1) (i.e., the element $t_{t_{\alpha_{2i}}(\alpha_{2i+1})}$ is in $G$ for $13\leq i\leq g-$ 1). Note that $\alpha_{2i-1}$ and $\alpha_{2i}$ are disjoint from $\alpha_{j}$ for $j\leq 2i-3$ or $2i+2\leq j$. Moreover, if $3\leq i$ then $\alpha_{2i-1},\alpha_{2i},\alpha_{2i+1}$ are disjoint from $\beta$. From these, we have
$f_{2i+1}(t_{\alpha_{2i-1}}(\alpha_{2i}))$ $=t_{\beta}t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}(t_{\alpha_{2i-1}}(\alpha_{2i}))$
$=t_{\beta}t_{\alpha_{2i+1}}(t_{\alpha_{2i-1}}(\alpha_{2i}))$
$=t_{\alpha_{2i+1}}(t_{\alpha_{2i-1}}(\alpha_{2i})).$
Since $\alpha_{2i-1},\alpha_{2i},\alpha_{2i+1}$ are disjoint from $\alpha_{3},\alpha_{5},\ldots,\alpha_{2i-3}$, we obtain
$F_{2i+1}(t_{\alpha_{2i-1}}(\alpha_{2i}))$ $=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}t_{\alpha_{2i-1}}f_{2i+1}(t_{\alpha_{2i-1}}(\alpha_{2i}))$
$=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}t_{\alpha_{2i-1}}(t_{\alpha_{2i+1}}(t_{\alpha_{2i-1}}(\alpha_{2i})))$
$=t_{\alpha_{2i-1}}(t_{\alpha_{2i+1}}(t_{\alpha_{2i-1}}(\alpha_{2i})))$
$=t_{\alpha_{2i-1}}t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}(\alpha_{2i}).$
From the above arguments, the braid relation $t_{\alpha_{2i-1}}t_{\alpha_{2i}}t_{\alpha_{2i-1}}=t_{\alpha_{2i}}t_{\alpha_{2i-1}}t_{\alpha_{2i}}$ and the relation $t_{\alpha_{2i}}t_{\alpha_{2i+1}}(\alpha_{2i})=\alpha_{2i+1}$, we have
$t_{t_{\alpha_{2i-1}}(\alpha_{2i})}F_{2i+1}(t_{\alpha_{2i-1}}(\alpha_{2i}))$ $=t_{t_{\alpha_{2i-1}}(\alpha_{2i})}(t_{\alpha_{2i-1}}t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}(\alpha_{2i}))$
$=t_{\alpha_{2i-1}}t_{\alpha_{2i}}t_{\alpha_{2i-1}}^{-1}t_{\alpha_{2i-1}}t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}(\alpha_{2i})$
$=t_{\alpha_{2i-1}}t_{\alpha_{2i}}t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}(\alpha_{2i})$
@@
$=t_{\alpha_{2i}}(\alpha_{2i+1}).$
This gives $t_{t_{\alpha_{2i}}(\alpha_{2i+1})}=t_{t_{t_{\alpha_{2i-1}}(\alpha_{2i})}F_{2i+1}(t_{\alpha_{2i-1}}(\alpha_{2i}))}$, and hence the element $t_{t_{\alpha_{2i}}(\alpha_{2i+1})}$ is in $G$ by $t_{t_{\alpha_{2i-1}}(\alpha_{2i})},t_{t_{\alpha_{2i-1}}(\alpha_{2i})}F_{2i+1}\in G$ (since $t_{t_{\alpha_{2i-1}}(\alpha_{2i})}$ and $F_{2i+1}$ are in $G$).
Second, we show the latter part of Lemma 3.4 (1) (i.e., the element $t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}$ is in $G$ for $13\leq i\leq g-$ 1). Note that $\alpha_{2i}$ and $\alpha_{2i+1}$ are disjoint from $\alpha_{j}$ for $2i+3\leq j$. Moreover, if $3\leq i$ then $\alpha_{2i-1},\alpha_{2i},\alpha_{2i+1},\alpha_{2i+2}$ are disjoint from $\beta$. From these and the relation $t_{\alpha_{2i+2}}^{-1}(\alpha_{2i+1})=t_{\alpha_{2i+1}}(\alpha_{2i+2})$, we have
$f_{2i-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))$ $=t_{\beta}t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))$
$=t_{\beta}t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))$
@@
Moreover, using the braid relation $t_{\alpha_{2i+1}}t_{\alpha_{2i}}t_{\alpha_{2i+1}}=t_{\alpha_{2i}}t_{\alpha_{2i+1}}t_{\alpha_{2i}}$, the following holds:
$t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}t_{\alpha_{2i+1}}t_{\alpha_{2i}}t_{\alpha_{2i+1}}$ $=t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}t_{\alpha_{2i}}t_{\alpha_{2i+1}}t_{\alpha_{2i}}$
$=t_{\alpha_{2i-1}}t_{\alpha_{2i+1}}t_{\alpha_{2i}}$
$=t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}t_{\alpha_{2i}}.$
Since $\alpha_{2i-1}$ and $\alpha_{2i}$ are disjoint from $\alpha_{2i+2}$, the above arguments give
$f_{2i-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))=t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}t_{\alpha_{2i}}(\alpha_{2i+2})=t_{\alpha_{2i+1}}(\alpha_{2i+2}),$
and moreover, since $\alpha_{2i+1}$ and $\alpha_{2i+2}$ are disjoint from $\alpha_{3},\alpha_{5},\ldots,\alpha_{2i-3}$, we have
$F_{2i-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))$ $=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}f_{2i-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))$
@@
$=t_{\alpha_{2i+1}}(\alpha_{2i+2}).$
This gives $t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}=t_{F_{2i-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))}$, and hence the element $t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}$ is in $G$ by $F_{2i-1},t_{t_{\alpha_{2i}}(\alpha_{2i+1})}\in G$.
Third, we show Lemma 3.4 (2) (i.e., the element $F_{2i+3}$ is in $G$ for $i\leq g-2$). Since $\alpha_{2i+1}$ and $\alpha_{2i+2}$ are disjoint from $\beta$ if $3\leq i$, by the relation $t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}t_{\alpha_{2i+1}}^{-1}=t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}^{-1}$ we obtain
@@
$=t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}t_{\beta}t_{\alpha_{2i+3}}t_{\alpha_{2i+4}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
@@
$=t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}^{-1}t_{\alpha_{2i+1}}\cdot f_{2i+3}.$
Since $\alpha_{2i+1}$ and $\alpha_{2i+2}$ are disjoint from $\alpha_{j}$ for $j\leq 2i-1$, we have
$G\ni F_{2i+1}$ $=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}t_{\alpha_{2i-1}}f_{2i+1}$
$=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}t_{\alpha_{2i-1}}\cdot t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}^{-1}t_{\alpha_{2i+1}}\cdot f_{2i+3}$
$=t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}^{-1}\cdot t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}t_{\alpha_{2i-1}}\cdot t_{\alpha_{2i+1}}\cdot f_{2i+3}$
$=t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}^{-1}F_{2i+3}.$
By $t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}\in G$ (i.e., Lemma 3.4 (1)), the element $F_{2i+3}$ is in $G$.
Finally, we show Lemma 3.4 (3) (i.e., the element $t_{\alpha_{2g}}$ is in $G$ for $i=g-1$). If $i=g-1$, then we have $F_{2i+1}=F_{2g-1}=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}f_{2g-1}\in G$. Since $\alpha_{2g-1}$ and $\alpha_{2g}$ are disjoint from $\beta,\alpha_{3},\alpha_{5},\ldots,\alpha_{2g-3}$, we see that the following holds:
$F_{2g-1}=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}f_{2g-1}$ $=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
@@
$=t_{t_{\alpha_{2g-1}(\alpha_{2g})}}\cdot t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}}.$
By $F_{2g-1}=t_{t_{\alpha_{2g-1}(\alpha_{2g})}}\cdot t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}}\in G$ and $t_{t_{\alpha_{2g-1}(\alpha_{2g})}}\in G$ (i.e., Lemma 3.4 (1)), we have $t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}}\in G$. Therefore, the element $t_{\alpha_{2g}}$ is in $G$ by $G\ni F_{2g-1}=(t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}})t_{\alpha_{2g}}^{-1}$, and the lemma follows. ∎
@@
We note that we have seen that the elements $t_{\alpha_{1}}$, $t_{\alpha_{2}}$, $t_{\alpha_{3}}$, $t_{t_{\alpha_{4}}(\alpha_{5})}$, $t_{t_{\alpha_{5}}(\alpha_{6})}$ and $f_{5}$ are contained in $G$ by Lemma 3.3.
Suppose that $g=3$. Since $\alpha_{5}$ and $\alpha_{6}$ are disjoint from $\beta$, we have
$G\ni f_{5}=t_{\beta}t_{\alpha_{5}}t_{\alpha_{6}}^{-1}=t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{5}}^{-1}t_{\beta}t_{\alpha_{5}}=t_{t_{\alpha_{5}}(\alpha_{6})}^{-1}\cdot t_{\beta}t_{\alpha_{5}},$
and hence the element $t_{\beta}t_{\alpha_{5}}$ is in $G$ by $t_{t_{\alpha_{5}}(\alpha_{6})}\in G$. Moreover, by $G\ni f_{5}=(t_{\beta}t_{\alpha_{5}})t_{\alpha_{6}}^{-1}$, we have $t_{\alpha_{6}}\in G$. Thus, since $t_{\alpha_{5}}(\alpha_{6})=t_{\alpha_{6}}^{-1}(\alpha_{5})$, we see that $t_{t_{\alpha_{5}}(\alpha_{6})}=t_{t_{\alpha_{6}^{-1}}(\alpha_{5})}(=t_{\alpha_{6}}^{-1}t_{\alpha_{5}}t_{\alpha_{6}}$), and hence the element $t_{\alpha_{5}}$ is in $G$ by $t_{t_{\alpha_{5}}(\alpha_{6})},t_{\alpha_{6}}\in G$. By applying this argument, with $t_{\alpha_{6}},t_{t_{\alpha_{5}}(\alpha_{6})}\in G$ replaced by $t_{\alpha_{5}},t_{t_{\alpha_{4}}(\alpha_{5})}\in G$, we find that the element $t_{\alpha_{4}}$ is in $G$. Therefore, the elements $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{6}}$ are contained in $G$, and hence we see that $t_{\beta}$ is in $G$ by $f_{1}=t_{\beta}t_{\alpha_{1}}t_{\alpha_{2}}^{-1}\cdots t_{\alpha_{5}}t_{\alpha_{6}}^{-1}\in G$. Since $\mathcal{M}_{g}$ is generated by $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{6}},t_{\beta}$ if $g=3$, we see that $G=\mathcal{M}_{g}$.
Suppose that $g\geq 4$. Since $\alpha_{5}$ and $\alpha_{6}$ are disjoint from $\beta$, we have
$G\ni f_{5}$ $=t_{\beta}t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{7}}t_{\alpha_{8}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
@@
$=t_{t_{\alpha_{5}}(\alpha_{6})}^{-1}t_{\alpha_{5}}f_{7}.$
By $t_{t_{\alpha_{5}}(\alpha_{6})}\in G$, the element $t_{\alpha_{5}}f_{7}$ is contained in $G$. Moreover, since $t_{\alpha_{3}}$ and $f_{5}$ are in $G$, we have $F_{5}=t_{\alpha_{3}}f_{5}\in G$ and $F_{7}=t_{\alpha_{3}}t_{\alpha_{5}}f_{7}\in G$. Therefore, the condition of Lemma 3.4 is satisfied for the case where $i=3$. By applying Lemma 3.4 repeatedly, we have $t_{t_{\alpha_{j}}(\alpha_{j+1})}\in G$ for $j=6,7,\ldots,2g-1$ and $t_{\alpha_{2g}}\in G$. Here, by $t_{\alpha_{2g-1}(\alpha_{2g})}=t_{\alpha_{2g}}^{-1}(\alpha_{2g-1})$, we have $t_{t_{\alpha_{2g-1}}(\alpha_{2g})}=t_{t_{\alpha_{2g}}^{-1}(\alpha_{2g-1})}(=t_{\alpha_{2g}}^{-1}t_{\alpha_{2g-1}}t_{\alpha_{2g}})$, and hence the element $t_{\alpha_{2g-1}}$ is in $G$ by $t_{t_{\alpha_{2g-1}}(\alpha_{2g})},t_{\alpha_{2g}}\in G$. By repeating this argument, we find that the elements $t_{\alpha_{4}},t_{\alpha_{5}},\ldots,t_{\alpha_{2g}}$ are contained in $G$. Moreover, from the definition of $f_{1}$, we have $t_{\beta}\in G$, and hence $G=\mathcal{M}_{g}$.
@@
Let $H^{\prime}$ be the subgroup of $\mathcal{H}_{g}$ generated by $\mathcal{R}$, which is defined in Lemma 2.7, and $r_{\pi}\mathcal{R}r_{\pi}^{-1}$, where $r_{\pi}$ is the rotation defined in Section 2. Then, $H^{\prime}=\mathcal{H}_{g}$ for $g\geq 2$, and $\mathcal{H}_{1}$ is generated by $t_{\alpha_{1}}$ and $t_{\alpha_{2}}$.
@@
Suppose that $g\geq 2$. From Lemma 2.7, we have $t_{\alpha_{2g+1}}^{-1}=\mathcal{R}^{2g-1}\in H^{\prime}$, and therefore $(t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}})^{g-1}=\mathcal{R}t_{\alpha_{2g+1}}\in H^{\prime}$. From these and the chain relation $(t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}})^{2g-1}=t_{\alpha_{2g+1}}^{2}$ (see the proof of Lemma 2.7), we obtain the following:
$t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}}=t_{\alpha_{2g+1}}^{2}(t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}})^{-2(g-1)}\in H^{\prime}.$
Therefore, by $H^{\prime}\ni(r_{\pi}\mathcal{R}r_{\pi}^{-1})^{2g-1}=r_{\pi}\mathcal{R}^{2g-1}r_{\pi}^{-1}=r_{\pi}t_{\alpha_{2g+1}}^{-1}r_{\pi}^{-1}=t_{r_{\pi}(\alpha_{2g+1})}^{-1}=t_{\alpha_{1}}^{-1}$, we have $t_{\alpha_{1}},t_{\alpha_{1}}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}}\in H^{\prime}$. It is easy to check that $t_{\alpha_{1}}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}}(\alpha_{j})=\alpha_{j+1}$ for $j=1,2,\ldots,2g-2$, and we have $t_{\alpha_{j+1}}=t_{f(\alpha_{j})}=ft_{\alpha_{j}}f^{-1}$, where $f=t_{\alpha_{1}}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}}$. From this and $t_{\alpha_{1}}\in H^{\prime}$, we obtain that the elements $t_{\alpha_{2}},t_{\alpha_{3}},\ldots,t_{\alpha_{2g-1}}$ are contained in $H^{\prime}$. Here, we see that
@@
$=((r_{\pi}t_{\alpha_{1}}r_{\pi}^{-1})^{2}\cdot r_{\pi}t_{\alpha_{2}}r_{\pi}^{-1}\cdot r_{\pi}t_{\alpha_{3}}r_{\pi}^{-1}\cdots r_{\pi}t_{\alpha_{2g-1}}r_{\pi}^{-1})^{g-1}r_{\pi}t_{\alpha_{2g+1}}^{-1}r_{\pi}^{-1}$

### v3

@@ 2. Preliminaries
First, we show the former part of Lemma 3.4 (1) (i.e., the element $t_{\gamma_{2i}}$ is in $G$ for $3\leq i\leq g-1$). Note that the following holds:
$f_{2i+1}(\gamma_{2i-1})$ $=t_{\beta}t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}(t_{\alpha_{2i-1}}(\alpha_{2i}))=t_{\beta}t_{\alpha_{2i+1}}(t_{\alpha_{2i-1}}(\alpha_{2i})).$
Therefore, by $3\leq i$, we have $f_{2i+1}(\gamma_{2i-1})=t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}(\alpha_{2i})$. This gives
$F_{2i+1}(\gamma_{2i-1})$ $=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}t_{\alpha_{2i-1}}(t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}(\alpha_{2i}))$
$=t_{\alpha_{2i-1}}(t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}(\alpha_{2i})).$
From the above arguments with $t_{\alpha_{2i}}t_{\alpha_{2i+1}}(\alpha_{2i})=\alpha_{2i+1}$, we have
$t_{\gamma_{2i-1}}F_{2i+1}(\gamma_{2i-1})$ $=t_{\alpha_{2i-1}}t_{\alpha_{2i}}t_{\alpha_{2i-1}}^{-1}t_{\alpha_{2i-1}}t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}(\alpha_{2i})$
@@
$=t_{\alpha_{2i}}(\alpha_{2i+1})(=\gamma_{2i}),$
and hence the element $t_{\gamma_{2i}}$ is in $G$.
Second, we show the latter part of Lemma 3.4 (1) (i.e., the element $t_{\gamma_{2i+1}}$ is in $G$ for $3\leq i\leq g-1$). By $3\leq i$, we have
$f_{2i-1}(\gamma_{2i})$ $=t_{\beta}t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))$
$=t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}t_{\alpha_{2i}}(\alpha_{2i+1})$
@@
Moreover, by $t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}t_{\alpha_{2i+1}}t_{\alpha_{2i}}t_{\alpha_{2i+1}}=t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}t_{\alpha_{2i}}t_{\alpha_{2i+1}}t_{\alpha_{2i}}=t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}t_{\alpha_{2i}}$, we obtain
$f_{2i-1}(\gamma_{2i})=t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}t_{\alpha_{2i}}(\alpha_{2i+2})=t_{\alpha_{2i+1}}(\alpha_{2i+2}),$
and moreover, we have
$F_{2i-1}(\gamma_{2i})$ $=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}f_{2i-1}(\gamma_{2i})$
@@
$=t_{\alpha_{2i+1}}(\alpha_{2i+2})(=\gamma_{2i+1}).$
Therefore, the element $t_{\gamma_{2i+1}}$ is in $G$.
Third, we show Lemma 3.4 (2) (i.e., the element $F_{2i+3}$ is in $G$ for $i\leq g-2$). By $3\leq i$, we obtain
@@
$=t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}t_{\alpha_{2i+1}}^{-1}\cdot t_{\alpha_{2i+1}}\cdot f_{2i+3}.$
Therefore, we have
$F_{2i+1}$ $=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}t_{\alpha_{2i-1}}\cdot t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}t_{\alpha_{2i+1}}^{-1}\cdot t_{\alpha_{2i+1}}\cdot f_{2i+3}$
$=t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}t_{\alpha_{2i+1}}^{-1}\cdot t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}t_{\alpha_{2i-1}}\cdot t_{\alpha_{2i+1}}\cdot f_{2i+3}$
$=t_{\gamma_{2i+1}}^{-1}F_{2i+3}.$
Since $t_{\gamma_{2i+1}}$ is in $G$, the element $F_{2i+3}$ is in $G$.
Finally, we show Lemma 3.4 (3) (i.e., the element $t_{\alpha_{2g}}$ is in $G$ for i=g1i=g-1). If i=g1i=g-1, then $F_{2i+1}=F_{2g-1}=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}f_{2g-1}$ is in $G$. Then, we see that the following holds:
$F_{2g-1}$ $=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
@@
$=t_{\gamma_{2g-1}}\cdot t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}}.$
This means that $t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}}$ is in $G$ since $t_{\gamma_{2i+1}}=t_{\gamma_{2g-1}}$ is in $G$. Therefore, the element $t_{\alpha_{2g}}$ is in $G$ by $F_{2g-1}=(t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}})t_{\alpha_{2g}}^{-1}$, and the lemma follows. ∎
@@
We note that we have seen that the elements $t_{\alpha_{1}}$, $t_{\alpha_{2}}$, $t_{\alpha_{3}}$, $t_{\gamma_{4}}$, $t_{\gamma_{5}}$ and $f_{5}$ are contained in $G$ by Lemma 3.3.
Suppose that $g=3$. Then, we have
$f_{5}=t_{\beta}t_{\alpha_{5}}t_{\alpha_{6}}^{-1}=t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{5}}^{-1}t_{\beta}t_{\alpha_{5}}=t_{\gamma_{5}}^{-1}\cdot t_{\beta}t_{\alpha_{5}},$
and hence the element $t_{\beta}t_{\alpha_{5}}$ is in $G$. Moreover, $t_{\alpha_{6}}$ is in $G$ by $f_{5}=(t_{\beta}t_{\alpha_{5}})t_{\alpha_{6}}^{-1}$, $t_{\alpha_{5}}$ is in $G$ by $\gamma_{5}=t_{\alpha_{5}}(\alpha_{6})=t_{\alpha_{6}}^{-1}(\alpha_{5})$, and $t_{\alpha_{4}}$ is in $G$ by $\gamma_{4}=t_{\alpha_{4}}(\alpha_{5})=t_{\alpha_{5}}^{-1}(\alpha_{4})$. Therefore, the elements $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{6}}$ are contained in $G$, and hence we see that $t_{\beta}$ is in $G$ by $f_{1}=t_{\beta}t_{\alpha_{1}}t_{\alpha_{2}}^{-1}\cdots t_{\alpha_{5}}t_{\alpha_{6}}^{-1}$. Since $\mathcal{M}_{g}$ is generated by $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{6}},t_{\beta}$ if $g=3$, we see that $G=\mathcal{M}_{g}$.
Suppose that $g\geq 4$. Then, we have
$f_{5}$ $=t_{\beta}t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{7}}t_{\alpha_{8}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
@@
$=t_{\gamma_{5}}^{-1}t_{\alpha_{5}}f_{7}.$
This means that the element $t_{\alpha_{5}}f_{7}$ is contained in $G$. Moreover, since $t_{\alpha_{3}}$ and $f_{5}$ are in $G$, $F_{5}=t_{\alpha_{3}}f_{5}$ and $F_{7}=t_{\alpha_{3}}t_{\alpha_{5}}f_{7}$ are in $G$. Therefore, the condition of Lemma 3.4 is satisfied for the case where $i=3$. By applying Lemma 3.4 repeatedly, we see that $t_{\gamma_{j}}$ is in $G$ for $j=6,7,\ldots,2g-1$ and $t_{\alpha_{2g}}\in G$, and hence the element $t_{\alpha_{2g-1}}$ is in $G$ by $\gamma_{2g-1}=t_{\alpha_{2g-1}(\alpha_{2g})}=t_{\alpha_{2g}}^{-1}(\alpha_{2g-1})$. By repeating this argument, we find that the elements $t_{\alpha_{4}},t_{\alpha_{5}},\ldots,t_{\alpha_{2g}}$ are contained in $G$. Moreover, from the definition of $f_{1}$, we have $t_{\beta}\in G$, and hence $G=\mathcal{M}_{g}$.
@@
Let $H^{\prime}$ be the subgroup of $\mathcal{H}_{g}$ generated by $\mathcal{R}$ and $r_{\pi}\mathcal{R}r_{\pi}^{-1}$, where $\mathcal{R}$ is defined in Lemma 2.7 and $r_{\pi}$ is the rotation defined in Section 2. Then, $H^{\prime}=\mathcal{H}_{g}$ for $g\geq 2$, and $\mathcal{H}_{1}$ is generated by $t_{\alpha_{1}}$ and $t_{\alpha_{2}}$.
@@
Suppose that $g\geq 2$. We set $f=t_{\alpha_{1}}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}}$. From Lemma 2.7, $t_{\alpha_{2g+1}}^{-1}=\mathcal{R}^{2g-1}$ is in $H^{\prime}$, and therefore $(t_{\alpha_{1}}f)^{g-1}=\mathcal{R}t_{\alpha_{2g+1}}$ is also in $H^{\prime}$. From these and the chain relation $(t_{\alpha_{1}}f)^{2g-1}=t_{\alpha_{2g+1}}^{2}$ (see the proof of Lemma 2.7), we see that $t_{\alpha_{1}}f=t_{\alpha_{2g+1}}^{2}(t_{\alpha_{1}}f)^{-2(g-1)}$ is contained in $H^{\prime}$. Therefore, by $(r_{\pi}\mathcal{R}r_{\pi}^{-1})^{2g-1}=r_{\pi}t_{\alpha_{2g+1}}^{-1}r_{\pi}^{-1}=t_{r_{\pi}(\alpha_{2g+1})}^{-1}=t_{\alpha_{1}}^{-1}$, the elements $t_{\alpha_{1}}$ and $f$ are in $H^{\prime}$. Since it is easy to check that $f(\alpha_{j})=\alpha_{j+1}$ for $j=1,2,\ldots,2g-2$, the elements $t_{\alpha_{2}},t_{\alpha_{3}},\ldots,t_{\alpha_{2g-1}}$ are contained in $H^{\prime}$. Here, we see that